# Can We Trust AI Agents? An Experimental Study Towards Trustworthy LLM-Based Multi-Agent Systems for AI Ethics

### v1

@@ 4 Results and Discussion
In this Section we present for each project description provided its corresponding results and discussion, finally, an overview discussion is presented. All project descriptions are derived from real AI incidents available at https://incidentdatabase.ai/. All data used in this experiment, including agent prompts, custom instructions, thematic analysis performed, are available at here.
@@ 4.1 Project Description 1
The first project description derived from this incident:
@@
It is worth highlighting that in the project description there is a need for referencing the EU AI Act, however, by using GPT-4o-mini, that was trained up until October 2023, this Act was not yet available, only being published on June 2024. In future work, we will introduce this document as a vector to be embedded and accessed by an open-source LLM. Moreover, according to David Lo [22], guaranteeing that AI4SE solution will comply with legislation is one way to enhance trustworthiness in AI4SE, since practitioners need to avoid conflict with laws.
@@
In regards to the ablation study, only 78 lines of text were obtained, while no code was produced. On the other hand, three themes were devised from the thematic analysis: Ethical AI Development, Technical Implementation and Fairness, and Compliance. It was not possible to generate the hierarchical clustering for any of the ablation studies due to small size of text data available. The quality between both studies is very disparate.
@@ 4.2 Project Description 2
The second project description derived from this incident:
@@
As the relationship between the thematic analysis in Table 2 and the hierarchical clustering produced in Figure 4, it is possible to highlight the class number 1 with theme number 2, Ethical Compliance and AI Responsibility. Moreover, class number 2 is closely related to theme Security and Collaboration, with words such as authentication and logger.
Similarly to the previous PD, no source code was produced in the ablation study. This yielded 78 lines of text and themes such as Technological Framework, Ethics and Compliance, User Experience and Support, Strategic Objective. While the study with the proposed prototype provides a more fruitful outcome, yielding almost 2000 lines of text, the ablation study provides a generic text that can only serve as study material.
@@ 4.3 Project Description 3
The third project description derived from this incident:
@@
This discussion follows in a similar fashion to the previous discussions. Raw data obtained vary from 1676 lines to 2122 lines of text data. The thematic analysis in Table 3 revolves around Ethical AI Development and Compliance and Technical Implementation, Optimization and Robustness. Class number 1 is related to theme number 1, however, the rest of the classes available contain words that are related to technical implementations, such as image_path, import, sklearn, sensitive_attribute, which are related to theme number 2. Interesting to note that sklearn is present in a class in Figure 5, even though it had already been recognized to have deprecated. This can highlight the fact that since its last update, in October 2023, and the date this study is being written, August 2024, it is hard for a AI4SE to follow the novelties of the technology while reinforcing the use of a single package for different tasks.
@@ 5 Threats to Validity
Different limitations appear from this study, including reliance on proprietary LLMs, which may change unpredictably. Reproducibility is a concern due to the non-deterministic outputs of LLMs. Since our methodology relies mainly on LLM to assess the outputs of the prototype, there is a potential risk of circular reasoning, where an AI is validating another AI’s trustworthiness. This could introduce bias, as both systems may share underlying flaws or limitations that go undetected. Moreover, thematic analysis and hierarchical clustering, though useful in organizing and analyzing text data, are both automated processes here, lacking human intervention. This raises concerns about the depth and interpretability of the identified themes. Without human input from ethicists or SE practitioners, there is a risk that the analysis may overlook nuanced or context-specific issues that only humans would detect, particularly when it comes to ethical concerns and practical challenges in SE workflows. To mitigate this, future work should incorporate human-in-the-loop methodologies where human stakeholders can review and validate the results produced by AI systems, offering insights and corrections that purely AI-driven approaches may miss. Furthermore, the evaluation of the source code was conducted at a surface level, with emphasis on ensuring that the sourc ecode executes without errors. However, no in-depth analysis was performed to assess whether the output aligns with AI ethics principles. Future work should incorporate a more rigorous assessment of the ethical implications of the generated source code, ensuring that it is not only functional but also ethically sound.
6 Final Remarks
@@
This study aimed to explore trustworthiness in AI4SE, specifically in the use of LLM for software engineering (LLM4SE), in the context of ethically aligned AI-based systems development. To the best of our knowledge, there are no studies in the literature that address practical AI ethics using LLM. Using the Design Science Research method, we identified the motivation and different techniques to improve trustworthiness in AI4SE, proposed a prototype - an LLM-based multi-agent system (LLM-BMAS) -, and evaluated it using thematic analysis, hierarchical clustering, ablation study, and source code execution. The techniques discovered to improve trustworthiness are a combination of agents, different roles (specialised agents), structured conversations and multiple rounds of debate. The prototype was tested with three distinct project descriptions derived from real AI incidents extracted from the AI Incident Database. For each project description, LLM-BMAS produced about 2000 lines of text, encompassing source code and documentation to implement the AI-based systems.
@@
Some of the codes produced through thematic analysis include: bias detection, transparency and accountability, user consent and GDPR compliance, fairness evaluation and compliance with the EU AI Act. In relation to hierarchical clustering, some of the words that depict ethics in AI include: ethical, bias, bias_type, evaluate_bias, bias_score. This highlight the ability of LLM-BMAS to generate outputs related to legal aspects of ethics in AI (i.e. GDPR and EU AI Act), in addition to theoretical and practical ethical issues, both in the documentation and in the source code, through codes in the thematic analysis (e.g., transparency, fairness, bias) and functions detected in the hierarchical clustering (e.g., evaluate_bias), respectively. Regarding the ablation study, for each project description it produced around 80 lines of output without source code, covering considerably less AI ethical issues. Therefore, this experiment underlines the potential of employing trustworthiness improvement techniques in LLMs to the development of more ethically aligned AI-based systems.
Nevertheless, there are some issues that hinder the synergy for practitioners [22]. Worth highlighting is the lack of practicality, from scraping the source code from the text - even more difficult when it comes to modules -, and installing packages and dealing with deprecated dependencies related to the LLM training date. While it is possible to improve trustworthiness and quality in the use of LLM4SE, there are still open issues to work on to make it more practical for practitioners.
@@
Whilst this experiment was conducted relying only on the internal knowledge of the Open AI LLM, it is important to note that providing the AI ethical agent with legislation (e.g., GDPR and EU AI Act) or AI ethical methods or guidelines (e.g., ECCOLA [39]) is a prospective future work that we intend to pursue. Moreover, in-depth analysis of the source code provided by the prototype is necessary to assess ethical alignment. With this study, we hope to further the understanding of the benefits of improving trustworthiness in LLM for the development of ethical AI-based systems. We aim to open source the prototype source code in further studies.
{credits}
6.0.1 Acknowledgements
@@
[…]
The authors have no competing interests to declare that are relevant to the content of this article.

### v2

@@ 4 Results and Discussion
In this Section, we present the results and discussion for each of the project descriptions used in the case study. Each project description is derived from real AI incidents available at https://incidentdatabase.ai/. The selected incidents focus on AI applications in recruitment, deepfake detection, and image classification. They were selected due to their high relevance to AI ethics and the need for compliance with regulatory standards. All data generated during the study – including agent prompts, custom instructions, and thematic analyses – are available in Zenodo.
@@ 4.1 Project Description 1
The first project description derived from Incident 37: Female Applicants Down-Ranked by Amazon Recruiting Tool:
@@
It is important to note that the project description references the EU AI Act, but the LLM used may not have had access to its full content at the time this study was conducted. In future work, we will introduce this document as a vector to be embedded and accessed by an open-source LLM. Moreover, according to David Lo [6], guaranteeing that AI4SE solution will comply with legislation is one way to enhance trustworthiness in AI4SE, since practitioners need to avoid conflict with laws.
@@
In regards to the baseline study, only 78 lines of text were obtained, while no code was produced. On the other hand, three themes were devised from the thematic analysis: Ethical AI Development, Technical Implementation and Fairness, and Compliance. It was not possible to generate the hierarchical clustering for any of the baseline studies due to small size of text data available. The quality between both studies is very disparate.
@@ 4.2 Project Description 2
As the relationship between the thematic analysis in Table 3 and the hierarchical clustering produced in Figure 4, it is possible to highlight the class number 1 with theme number 2, Ethical Compliance and AI Responsibility. Moreover, class number 2 is closely related to theme Security and Collaboration, with words such as authentication and logger. Similarly to the previous PD, no source code was produced in the baseline study. This yielded 78 lines of text and themes such as Technological Framework, Ethics and Compliance, User Experience and Support, Strategic Objective. While the study with the proposed prototype provides a more fruitful outcome, yielding almost 2,000 lines of text, the baseline study provides a generic text that can only serve as study material.
@@ 4.3 Project Description 3
The third project description derived from Incident 16: Images of Black People Labeled as Gorillas:
@@
This discussion follows in a similar fashion to the previous discussions. Raw data obtained vary from 1676 lines to 2122 lines of text data. The thematic analysis in Table 4 revolves around Ethical AI Development and Compliance and Technical Implementation, Optimization and Robustness. Class number 1 is related to theme number 1, however, the rest of the classes available contain words that are related to technical implementations, such as image_path, import, sklearn, sensitive_attribute, which are related to theme number 2. Interesting to note that sklearn is present in a class in Figure 5, even though it had already been recognized to have deprecated. This highlights the challenge of relying on static LLM knowledge in a rapidly evolving software landscape, where tools and dependencies can quickly become outdated.
@@ 5 Threats to Validity
This study has several limitations that may affect the validity and generalizability of its findings. First, the prototype relies on a proprietary LLM by OpenAI, which may evolve over time in undocumented ways. This creates a reproducibility challenge, as future versions of the model may behave differently, even under identical conditions. Secondly, the evaluation relies on an LLM to perform a thematic analysis of the output, which could raise questions regarding its validity and objectivity. Although recent studies (e.g. [10]; [36]) have demonstrated the potential of LLMs to support, and even enhance, qualitative analysis by identifying patterns that human coders might overlook, this approach remains methodologically debated. In our study, we did not conduct a parallel human-led thematic analysis, which limits our ability to triangulate or verify the consistency of the themes generated. This increases the risk that the LLM-based analysis will reflect the model’s own limitations or biases. Future work should either involve expert human analysts to validate the AI-generated codes and themes or adopt a hybrid human-in-the-loop approach to enhance analytical rigour. Third, the thematic analysis and hierarchical clustering processes were fully automated and lacked human validation. While these methods offer efficiency and consistency, they may miss subtle, context-sensitive insights that would be more easily detected by domain experts, particularly in the ethical dimensions of software engineering.
Fourth, although the prototype generates source code, our execution tests were limited to confirming whether the code runs, rather than assessing its functionality, correctness, or ethical alignment. This leaves open questions about the practical reliability and appropriateness of the generated software. Finally, the study was limited to three case scenarios. While these were selected for their ethical relevance, the small sample size restricts the scope of generalizability. Future work should expand on this with broader case coverage, incorporate human-in-the-loop validation, and include a more rigorous assessment of the ethical dimensions of the generated artifacts. It is also important to note that the evaluation was conducted using a case study approach. Our objective was not to produce generalisable results, but rather to explore and gain in-depth insights into how trustworthiness-enhancing techniques in LLM-based systems can support the development of ethically aligned AI. Case studies are particularly valuable for examining complex, real-world phenomena in context, and for uncovering practical challenges and opportunities that may not emerge in more controlled or large-scale experiments [40].
6 Conclusion
@@
This study aimed to explore trustworthiness in AI4SE, specifically in the use of LLM for software engineering (LLM4SE), in the context of ethically aligned AI-based systems development. To the best of our knowledge, there are no studies in the literature that address practical AI ethics using LLM. Our Research Question is: To what extent can trustworthiness-enhancing techniques in LLM-based systems support the development of ethically aligned AI software. In order to provide an answer, we used the Design Science Research method, we identified the motivation and different trustworthiness-enhancing techniques in LLM4SE, proposed a prototype - an LLM-based multi-agent system (LLM-MAS) -, and evaluated it through three case studies. The techniques discovered to improve trustworthiness are (1) multi-agent collaboration, (2) specialized roles, (3) structured communication, and (4) multiple rounds of debate. Each case study had a project description derived from real AI incidents extracted from the AI Incident Database in recruitment, deepfake detection, and image classification. For each project description, LLM-MAS produced about 2,000 lines of text, encompassing source code and documentation to implement the AI-based systems.
@@
Some of the codes produced through thematic analysis include: bias detection, transparency and accountability, user consent and GDPR compliance, fairness evaluation and compliance with the EU AI Act. In relation to hierarchical clustering, some of the words that depict ethics in AI include: ethical, bias, bias_type, evaluate_bias, bias_score. This highlight the ability of LLM-MAS to generate outputs related to legal aspects of ethics in AI (i.e. GDPR and EU AI Act), in addition to theoretical and practical ethical issues, both in the documentation and in the source code, through codes in the thematic analysis (e.g., transparency, fairness, bias) and functions detected in the hierarchical clustering (e.g., evaluate_bias), respectively. Regarding the baseline study, for each project description it produced around 80 lines of output without source code, covering considerably less AI ethical issues. Therefore, this study underlines the potential of employing trustworthiness-enhancing techniques in LLMs to the development of more ethically aligned AI-based systems.
@@
Nevertheless, there are some issues that hinder the synergy for practitioners [6]. Worth highlighting is the lack of practicality, from scraping the source code from the text - even more difficult when it comes to modules -, and installing packages and dealing with deprecated dependencies stemming from the LLM training date. While it is possible to improve trustworthiness and quality in the use of LLM4SE, there are still open issues to work on to make it more practical for practitioners. Whilst this experiment was conducted relying only on the internal knowledge of the Open AI LLM, it is important to note that providing the AI ethical agent with legislation (e.g., GDPR and EU AI Act) or AI ethical methods or guidelines (e.g., ECCOLA [1]) is a prospective future work that we intend to pursue. Moreover, in-depth analysis of the source code provided by the prototype is necessary to assess ethical alignment. With this study, we hope to further the understanding of the benefits of trustworthiness-enhancing techniques in LLM for the development of ethical AI-based systems. We aim to open source the prototype source code in further studies.
Acknowledgements.